OpenStudy (anonymous):

This is a literal equation and I'm kind of confused about this one: Solve 4x + t/r = s for x

OpenStudy (tkhunny):

Is it \(\dfrac{4x + t}{r} = s\;or\;4x + \dfrac{t}{r} = s\) You have written the latter.